A baseball of mass 0.239 kg sits on a tee. A kid's bat hits the ball and maintains contact for 0.005 seconds. The result is that the ball leaves the tee with a velocity of 11.03 m/s, traveling exactly horizontally.



Find the impulse given to the ball.
1
Expert's answer
2020-04-16T09:11:37-0400

Impulse given to the ball"=" change in momentum of the ball

So, impulse given"=m(v_2-v_1)=0.239\times(11.03-0)=2.63617Ns"